# Finite element approximation of the stationary Navier-Stokes problem with non-smooth data

### v1

@@ Abstract
The aim of this work is to analyze the finite element approximation of the stationary Navier-Stokes equations with non-smooth Dirichlet boundary data. The discrete approximation is obtained by considering the Navier-Stokes system with a regularized boundary data. Based on the existence of the very weak solution, for the Navier–Stokes system with $L^{2}$ boundary data, and a suitable descomposition of this solution, we obtain a priori error estimates between the approximation of the Navier-Stokes system with non-smooth data and the finite element solution of the associated regularized problem. These estimates allow us to conclude that our approach converges with optimal order.
@@
2020 MSC:
65N15 , 65N30
[…]
[label]organization=Departamento de Matemática, Facultad de Ciencias Exactas y Naturales, Universidad de Buenos Aires, IMAS - Conicet, city=Buenos Aires, postcode=1428, country=Argentina
@@ 1 Introduction
In this work we analyze the finite element approximation of the stationary Navier-Stokes problem with non-smooth Dirichlet data. One motivation to study this problem comes from the so-called cavity flow problem, which is a classic benchmark for testing numerical methods, in particular finite element methods. A natural way to obtain the discrete solution of the problem is the regularization of the boundary data combined with the weak solution of the regularized problem. Recent studies have focused on solving the Poisson problem under $L^{2}$ boundary conditions [1, 2]. More recently, [3, 4] investigates this issue for the Stokes problem and [5] addressed it for linear elliptic problems. The nonlinearity of the Navier-Stokes equations introduces an additional difficulty, which prevents direct generalization of these results.
First, we introduce the notation that will be used throughout the paper. Let $s\geq 0$ be a real number, $1\leq p\leq\infty$ and $D\subset\mathbb{R}^{2}$ be an open bounded set, we denote $W^{s,p}(D)$ to the standard space of Sobolev in $D$, with $\|\cdot\|_{W^{s,p}(D)}$ and $|\cdot|_{W^{s,p}(D)}$ its norm and seminorm respectively (see, for instance, [6]). The space $W^{s,p}_{0}(D)$ is defined as the closure of $C^{\infty}$ functions with compact support in $D$, i.e, the clousure of $C^{\infty}_{0}(\bar{D})$ in $W^{s,p}(D)$. For $p=2$ we write $H^{s}(D)=W^{s,2}(D)$ and $H^{s}_{0}(D)=W_{0}^{s,2}(D)$.
$W^{-s,q}(D)$ is defined as the dual space of $W_{0}^{s,p}(D)$ with $\frac{1}{p}+\frac{1}{q}=1$. In $H^{-1}(D)$ we define the norm
@@
On the other hand we will say that a function $f(x)=O(g(x))$ if there exists a constant $C>0$ such that $f(x)\leq Cg(x)$.
As usual, we use bold characters to denote vector-valued functions and their corresponding functional spaces.
@@
where $\mathbf{n}$ denotes the outward unit vector normal to the boundary.
The problem is find $(\textit{{u}},p)$ such that:
@@
If we assume $\mathbf{g}\in\mathbf{H}^{1/2}(\Gamma)$, the typical approach to solving the problem using finite elements is to multiply the equations by test functions and integrate by parts, leading to a suitable variational form in the standard spaces $\mathbf{H}^{1}(\Omega)$ for the velocity u and $L_{0}^{2}(\Omega)$ for the pressure $p$ (see, for example, [7, 8]). The challenge arises when $\mathbf{g}\notin\mathbf{H}^{1/2}(\Gamma)$, as in this case the solution $\textit{{u}}\notin\mathbf{H}^{1}(\Omega)$, and thus the standard theory cannot be applied.
@@
Although a similar approach is given in [3] for the Stokes problem, the nonlinearity of the Navier-Stokes equations introduces extra difficulties and requires a different treatment. To analyze the approximation, we first study the existence of solution to the problem (2) for a non-regular boundary data through the concept of a very weak solution. Then, we use an useful descomposition of the solution given in [9].
This descomposition allows us to show that the difference between the very weak solution of (2) and the solution of (3) can be controlled by the approximation error between the non-regular data $\mathbf{g}$ and the regularized data $\mathbf{g}_{\varepsilon}$. Then, we approximate the classical solution of (3) by finite elements and show the a priori error estimates under approriate assumptions of the finite element spaces and the meshes.
@@ 2 Preliminaries
For $\textit{{u}},\textit{{v}},\textit{{w}}\in\mathbf{H}^{1}(\Omega)$ and $p\in L_{0}^{2}(\Omega)$ we define:
$a(\textit{{u}},\textit{{v}})=\nu\int_{\Omega}\nabla\textit{{u}}:\nabla\textit{{v}},\quad b(\textit{{v}},q)=\int_{\Omega}q\textup{div}\ \textit{{v}},\,{\mbox{ and }}\,c(\textit{{u}},\textit{{v}},\textit{{w}})=\int_{\Omega}(\textit{{u}}\cdot\nabla)\textit{{v}}\cdot\textit{{w}}.$
Throughout the paper, we will denote by $C$ a generic positive constant, not necessarily the same at each occurrence.
Now we present some useful results [10, 11].
@@
Item (a) is shown in [11, lemma 1.1, Chapter II] where it is also shown that, for $\textit{{u}},\textit{{v}},\textit{{w}}\in\mathbf{H}^{1}_{0}(\Omega)$, there exists $C>0$ such that $|c(\textit{{u}},\textit{{v}},\textit{{w}})|\leq C\|\textit{{u}}\|_{H^{1}(\Omega)}\|\textit{{v}}\|_{H^{1}(\Omega)}\|\textit{{w}}\|_{H^{1}(\Omega)}$. Consequently, by applying Poincaré’s inequality, we derive (b). Items (c) and (d) are proven in [11, lemma 1.3, Chapter II]. ∎
@@
Given a function $\textit{{v}}_{0}\in\mathbf{H}^{1/2}(\Gamma)$ satisfying the condition (1) (i.e., $\int_{\Gamma}\textit{{v}}_{0}\cdot\mathbf{n}=0$), there exists, for any $\varepsilon>0$, a function $\Psi=\Psi(\varepsilon)\in\mathbf{H}^{1}(\Omega)$ such that
$\textup{div}\ \Psi=0,\qquad\Psi|_{\Gamma}=\textit{{v}}_{0}$
@@
Let $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$ and $\textit{{v}}_{0}\in\mathbf{H}^{1/2}(\Gamma)$ satisfy (1). Let us consider the problem of finding $(\textit{{v}},q)$ such that
$\left\{\begin{array}[]{rcc}-\nu\Delta\textit{{v}}+(\textit{{v}}\cdot\nabla)\textit{{v}}+\nabla q&=\mathbf{f}&\textup{in }\,\Omega\\ \textup{div}\ \textit{{v}}&=0&\textup{in }\,\Omega\\ \textit{{v}}&=\textit{{v}}_{0}&\textup{on }\,\Gamma\end{array}\right$. (4)
Given $\varepsilon>0$, if
@@
The proof follows the ideas given in [10, Theorem 2.3, Chapter IV] and [11, Theorems 1.5 and 1.6, Chapter II].
A variational form of problem (4) consists in finding a pair $(\textit{{v}},q)\in\mathbf{H}^{1}(\Omega)\times L_{0}^{2}(\Omega)$ solution of the equations
$\left\{\begin{array}[]{rll}a(\textit{{v}},\textit{{w}})+c(\textit{{v}},\textit{{v}},\textit{{w}})-b(\textit{{w}},q)&=\int_{\Omega}\mathbf{f}\cdot\textit{{w}}&\forall\textit{{w}}\in\mathbf{H}^{1}_{0}(\Omega)\\ \textup{div}\ \textit{{v}}&=0&\textup{ in }\,\Omega\\ \textit{{v}}&=\textit{{v}}_{0}&\textup{ on }\,\Gamma\end{array}\right$. (5)
Let $\Psi\in\mathbf{H}^{1}(\Omega)$ be the extension function of $\textit{{v}}_{0}$, with $\textup{div}\ \Psi=0$, given by the previous lemma and we define $\mathbf{V}=\{\textit{{w}}\in\mathbf{H}^{1}_{0}:\textup{div}\ \textit{{w}}=0\}$. We set $\textit{{v}}=\hat{\textit{{v}}}+\Psi$, with $\hat{\textit{{v}}}\in\mathbf{H}^{1}_{0}(\Omega)$. Since,
$a(\textit{{v}},\textit{{w}})+c(\textit{{v}},\textit{{v}},\textit{{w}})=a(\hat{\textit{{v}}},\textit{{w}})+a(\Psi,\textit{{w}})+c(\hat{\textit{{v}}},\hat{\textit{{v}}},\textit{{w}})+c(\Psi,\hat{\textit{{v}}},\textit{{w}})+c(\hat{\textit{{v}}},\Psi,\textit{{w}})+c(\Psi,\Psi,\textit{{w}}).$
The problem (5) is equivalent to:
Find a pair $(\hat{\textit{{v}}},q)\in\mathbf{V}\times L^{2}_{0}(\Omega)$ such that
$\tilde{a}(\hat{\textit{{v}}},\hat{\textit{{v}}},\textit{{w}})=\int_{\Omega}\hat{\mathbf{f}}\cdot\textit{{w}}\quad\forall\textit{{w}}\in\mathbf{V}$ (6)
@@
The proof of existence follows as in Theorem IV.2.1 of [10], noting that for all $\tilde{\textit{{w}}},\textit{{w}}\in\mathbf{V}$
@@
where we use item c) of lemma 2.1, for all $\textit{{w}},\tilde{\textit{{w}}}\in\mathbf{V}$. By Lemma 2.2, if we take $\varepsilon<\nu$, we have
$\tilde{a}(\tilde{\textit{{w}}},\textit{{w}},\textit{{w}})\geq(\nu-\varepsilon)|\textit{{w}}|^{2}_{H^{1}(\Omega)},\quad\forall\tilde{\textit{{w}}},\textit{{w}}\in\mathbf{V}.$
@@
Now, suppose that $\textit{{v}}_{1}$ and $\textit{{v}}_{2}$ are two solutions of (5). Then, $\hat{\textit{{v}}}_{1}=\textit{{v}}_{1}-\Psi$ and $\hat{\textit{{v}}}_{2}=\textit{{v}}_{2}-\Psi$ satisfy (6) and so for all $\textit{{w}}\in\mathbf{V}$
@@
and since $\left(\frac{\mathcal{C}_{NL}\|\hat{\mathbf{f}}\|_{H^{-1}(\Omega)}}{(\nu-\varepsilon)^{2}}\right)<1$ we conclude that $\hat{\textit{{v}}}_{1}-\hat{\textit{{v}}}_{2}=0.$
@@ 2.1 Existence of very weak solution
For linear problems, like the Poisson problem [1] or the Stokes problem [4, 3] with Dirichlet condition in $L^{2}(\Gamma)$, the corresponding very weak solution u belongs to the space $\mathbf{L}^{2}(\Omega)$. However, in our problem if $\textit{{u}}\in\mathbf{L}^{2}(\Omega)$ to ensure that
@@
We note that, if we consider the solution of the variational problem with a regular Dirichlet data (for example, in $\mathbf{H}^{1/2}(\Gamma)$), in which case $(\textit{{u}},p)\in\mathbf{H}^{1}(\Omega)\times L^{2}_{0}(\Omega)$, then u is also a very weak solution. Indeed, from 1.4.4.5 in [12], we have that $W^{2,4/3}(\Omega)\subset W^{1,4}(\Omega)$, and since $L^{4}(\Omega)\subset L^{2}(\Omega)$, we can conclude that if $\phi\in\mathbf{X}$ then $\phi\in\mathbf{H}^{1}(\Omega)$. On the other hand, since $W^{1,4/3}(\Omega)\subset L^{4}(\Omega)$ (see 1.4.4.5 in [12] again), we can also deduce that if $q\in M$, then $q\in L^{2}_{0}(\Omega)$.
It is important to note that $\mathbf{H}^{1}(\Omega)\hookrightarrow\mathbf{L}^{4}(\Omega)$ [6, 12]. Furthermore, given $\textit{{v}}\in\mathbf{H}^{1}(\Omega)$ there exists $C_{I}>0$ such that
@@
$\left\{\begin{array}[]{rcc}-\nu\Delta\textit{{w}}+\nabla\xi&=\mathbf{F}&\textup{in }\,\Omega\\ \textup{div}\ \textit{{w}}&=0&\textup{in }\,\Omega\\ \textit{{w}}&=0&\textup{on }\,\Gamma\end{array}\right$.
has a unique solution $(\textit{{w}},\xi)\in\mathbf{W}^{2,4/3}(\Omega)\times W^{1,4/3}(\Omega)\setminus\mathbb{R}$ and there exists a constant $C_{S}$ such that:
$\|\textit{{w}}\|_{W^{2,4/3}(\Omega)}+\|\xi\|_{W^{1,4/3}(\Omega)}\leq C_{S}\|F\|_{L^{4/3}(\Omega)}$
@@
As stated in [9], if $\Omega$ is a bounded $C^{1,1}$ domain, the Assumption 1 is satisfied according to Theorem 3 of [13].
The assumption also holds if $\Omega$ is a convex polygon (see [10, Theorem 5.4, Reamark 5.6] and [14]).
The first important result to consider from Marušić-Paloka work [9] is the existence and uniqueness of the very weak solution u for the case $\mathbf{f}=0$.
@@
Assume that $\Omega\subset\mathbb{R}^{2}$ satisfies Assumption 1,
$\mathbf{g}\in\mathbf{L}^{2}(\Gamma)$ satisfy (1), $\mathbf{f}=0$ and $\mathbf{L}_{\Gamma}=\{\textit{{w}}\in\mathbf{L}^{4}(\Omega),\quad\textup{div}\ \textit{{w}}=0,\quad\textit{{w}}\cdot\mathbf{n}\in L^{2}(\Gamma)\}$. If $\|\mathbf{g}\|_{L^{2}(\Gamma)}$ is small enough, then there exists a unique very weak solution $\textit{{u}}\in\mathbf{L}_{\Gamma}$ of (2). Furthermore, there exists a constant $C_{1}=C_{1}(\Omega)>0$ such that
@@
The result follows from [9, Theorem 4]. Indeed, if the assumption 1 is satisfied, the estimate provided in [9, Lemma 2] can be obtained. Consequently, as the author note in their Remark 3, all the results from that work are applicable. ∎

### v2

@@ Abstract
The aim of this work is to analyze the finite element approximation of the two-dimensional stationary Navier-Stokes equations with non-smooth Dirichlet boundary data. The discrete approximation is obtained by considering the Navier-Stokes system with a regularized boundary data. Based on the existence of the very weak solution, for the Navier–Stokes system with $L^{2}$ boundary data, and a suitable decomposition of this solution, we obtain a priori error estimates between the approximation of the Navier-Stokes system with non-smooth data and the finite element solution of the associated regularized problem. These estimates allow us to conclude that our approach converges with optimal order.
@@
pacs:
MSC Classification]65N15, 65N30
@@ 1 Introduction
The numerical approximation of incompressible flows with non-smooth boundary data is a recurring challenge in computational fluid dynamics. In many physically relevant situations, such as internal flows with sharp corners or driven boundary conditions, the prescribed boundary velocity lacks the regularity typically assumed in theoretical and numerical analyses. A prototypical example is the lid-driven cavity flow, which remains a canonical benchmark for numerical methods for the Stokes and Navier–Stokes equations.
In this work, we analyze the finite element approximation of the stationary Navier-Stokes problem with non-smooth Dirichlet data. When Dirichlet boundary data are sufficiently regular, the stationary Navier–Stokes equations admit weak solutions in the standard Sobolev spaces, and their finite element approximation is well understood (see, for example, [wang2016, gunzburger1983, fix1983, zheng] and the references therein). However, if the boundary data belong only to $L^{2}$, the classical variational formulation breaks down. This lack of regularity prevents a direct application of standard finite element theory and raises fundamental questions regarding the convergence and accuracy of numerical approximations.
Our approach is based on a regularization of the boundary data combined with the notion of very weak solutions for the original problem. More precisely, the non-regular boundary data are approximated by a family of smooth functions, leading to a regularized Navier–Stokes problem that admits a classical weak solution. The finite element method is then applied to this regularized problem.
In recent years, several works have addressed elliptic and linear fluid problems with low-regularity boundary data. Finite element approximations for the Poisson problem under $L^{2}$ boundary conditions were studied in [apel2016, apel2017]. The error estimation for the two-dimensional Stokes driven cavity problem was analyzed in [CAI2009]. More recently, [duran2020, hamouda2017] investigated this issue for the Stokes problem, and [CAI2023] addressed it for linear elliptic problems.
Extending these ideas to the Navier–Stokes equations is non-trivial due to the presence of the nonlinear convective term, which prevents a straightforward generalization of the linear theory.
First, we introduce the notation that will be used throughout the paper. Let $s\geq 0$ be a real number, $1\leq p\leq\infty$ and $D\subset\mathbb{R}^{2}$ be an open bounded set, we denote $W^{s,p}(D)$ to the standard space of Sobolev in $D$, with $\|\cdot\|_{W^{s,p}(D)}$ and $|\cdot|_{W^{s,p}(D)}$ its norm and seminorm respectively (for more details, see, for instance, [Adams1975]). The space $W^{s,p}_{0}(D)$ is defined as the closure of $C^{\infty}$ functions with compact support in $D$, i.e, the clousure of $C^{\infty}_{0}(\bar{D})$ in $W^{s,p}(D)$. For $p=2$ we write $H^{s}(D)=W^{s,2}(D)$ and $H^{s}_{0}(D)=W_{0}^{s,2}(D)$. $W^{-s,q}(D)$ is defined as the dual space of $W_{0}^{s,p}(D)$ with $\frac{1}{p}+\frac{1}{q}=1$. In $H^{-1}(D)$ we define the norm
@@
On the other hand we will say that a function $f(x)=O(g(x))$ if there exists a constant $C>0$ such that $f(x)\leq Cg(x)$. As usual, we use bold characters to denote vector-valued functions and their corresponding functional spaces.
@@
where $\mathbf{n}$ denotes the outward unit vector normal to the boundary. The stationary Navier-Stokes problem is find $(\textit{{u}},p)$ such that:
@@
If we assume $\mathbf{g}\in\mathbf{H}^{1/2}(\Gamma)$, the typical approach to solving the problem using finite elements is to multiply the equations by test functions and integrate by parts, leading to a suitable variational form in the standard spaces $\mathbf{H}^{1}(\Omega)$ for the velocity u and $L_{0}^{2}(\Omega)$ for the pressure $p$ (see, for example, [wang2016, fix1983, gunzburger1983]). The challenge arises when $\mathbf{g}\notin\mathbf{H}^{1/2}(\Gamma)$, as in this case the solution $\textit{{u}}\notin\mathbf{H}^{1}(\Omega)$, and thus the standard theory cannot be applied.
@@
Although a similar approach is given in [CAI2009, duran2020, hamouda2017] for the Stokes problem, the presence of the nonlinear convective term in the Navier-Stokes equations introduces extra difficulties and requires a different treatment. To analyze the approximation, we first study the existence of solution to the problem (2) for a non-regular boundary data through the concept of a very weak solution. To this end, we employ an useful decomposition of the solution given in [marusic2000]. This decomposition allows us to show that the difference between the very weak solution of (2) and the solution of (3) can be controlled by the approximation error between the non-regular data $\mathbf{g}$ and the regularized data $\mathbf{g}_{\varepsilon}$. Then, we approximate the classical solution of (3) by finite elements and show the a priori error estimates under approriate assumptions of the finite element spaces and the meshes.
@@ 2 Preliminaries
For our analysis, we first consider the following problem with regular data:
$\left\{\begin{array}[]{rcc}-\nu\Delta\textit{{v}}+(\textit{{v}}\cdot\nabla)\textit{{v}}+\nabla q&=\mathbf{f}&\textup{in }\,\Omega\\ \textup{div}\ \textit{{v}}&=0&\textup{in }\,\Omega\\ \textit{{v}}&=\textit{{v}}_{D}&\textup{on }\,\Gamma\end{array}\right$. (4)
with $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$ and $\textit{{v}}_{D}\in\mathbf{H}^{1/2}(\Gamma)$.
A variational form of problem (4) consists in finding a pair $(\textit{{v}},q)\in\mathbf{H}^{1}(\Omega)\times L_{0}^{2}(\Omega)$ solution of the equations, such that, $\textit{{v}}|_{\Gamma}=\textit{{v}}_{D}$ and
[…]
with
$a(\textit{{v}},\textit{{w}})=\nu\int_{\Omega}\nabla\textit{{v}}:\nabla\textit{{w}},\quad b(\textit{{w}},q)=\int_{\Omega}q\textup{div}\ \textit{{w}},\,{\mbox{ and }}\,c(\textit{{u}},\textit{{v}},\textit{{w}})=\int_{\Omega}(\textit{{u}}\cdot\nabla)\textit{{v}}\cdot\textit{{w}}.$
Throughout the paper, we will denote by $C$ a generic positive constant, not necessarily the same at each occurrence. Now we present some useful results [girault1986, temam1977].
@@
Item (a) is shown in [temam1977, lemma 1.1, Chapter II] where it is also shown that, for $\textit{{u}},\textit{{v}},\textit{{w}}\in\mathbf{H}^{1}_{0}(\Omega)$, there exists $C>0$ such that $|c(\textit{{u}},\textit{{v}},\textit{{w}})|\leq C\|\textit{{u}}\|_{H^{1}(\Omega)}\|\textit{{v}}\|_{H^{1}(\Omega)}\|\textit{{w}}\|_{H^{1}(\Omega)}$. Consequently, by applying Poincaré’s inequality, we derive (b). Items (c) and (d) are proven in [temam1977, lemma 1.3, Chapter II]. ∎
@@
Given a function $\textit{{v}}_{D}\in\mathbf{H}^{1/2}(\Gamma)$ satisfying the condition (1) (i.e., $\int_{\Gamma}\textit{{v}}_{D}\cdot\mathbf{n}=0$), there exists, for any $\varepsilon>0$, a function $\Psi=\Psi(\varepsilon)\in\mathbf{H}^{1}(\Omega)$ such that
$\textup{div}\ \Psi=0,\qquad\Psi|_{\Gamma}=\textit{{v}}_{D}$
@@
Let $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$ and $\textit{{v}}_{D}\in\mathbf{H}^{1/2}(\Gamma)$ satisfy (1). Given $\varepsilon>0$, if
@@
The proof follows the ideas given in [girault1986, Theorem 2.3, Chapter IV] and [temam1977, Theorems 1.5 and 1.6, Chapter II]. Let $\Psi\in\mathbf{H}^{1}(\Omega)$ be the extension function of $\textit{{v}}_{D}$, with $\textup{div}\ \Psi=0$, given by the previous lemma. We set $\textit{{v}}=\hat{\textit{{v}}}+\Psi$, with $\hat{\textit{{v}}}\in\mathbf{H}^{1}_{0}(\Omega)$. So,
$a(\textit{{v}},\textit{{w}})+c(\textit{{v}},\textit{{v}},\textit{{w}})$ $=a(\hat{\textit{{v}}},\textit{{w}})+a(\Psi,\textit{{w}})+c(\hat{\textit{{v}}},\hat{\textit{{v}}},\textit{{w}})$
$+c(\Psi,\hat{\textit{{v}}},\textit{{w}})+c(\hat{\textit{{v}}},\Psi,\textit{{w}})+c(\Psi,\Psi,\textit{{w}})$
Since the bilinear form $b(\cdot,\cdot)$ satisfies the inf–sup condition (see [girault1986, Theorem 1.4, Chapter IV]) problem (5) is equivalent to find $(\hat{\textit{{v}}},q)\in\mathbf{W}\times L^{2}_{0}(\Omega)$, with $\mathbf{W}=\{\textit{{w}}\in\mathbf{H}^{1}_{0}:\textup{div}\ \textit{{w}}=0\}$, such that
$\tilde{a}(\hat{\textit{{v}}},\hat{\textit{{v}}},\textit{{w}})=\int_{\Omega}\hat{\mathbf{f}}\cdot\textit{{w}}\quad\forall\textit{{w}}\in\mathbf{W},$ (6)
@@
The proof of existence follows as in Theorem IV.2.1 of [girault1986], noting that for all $\tilde{\textit{{w}}},\textit{{w}}\in\mathbf{W}$
@@
where we use item (c) of lemma 2.1, for all $\textit{{w}},\tilde{\textit{{w}}}\in\mathbf{W}$. By Lemma 2.2, if we take $\varepsilon<\nu$, we have
$\tilde{a}(\tilde{\textit{{w}}},\textit{{w}},\textit{{w}})\geq(\nu-\varepsilon)|\textit{{w}}|^{2}_{H^{1}(\Omega)},\quad\forall\tilde{\textit{{w}}},\textit{{w}}\in\mathbf{W}.$
@@
Now, suppose that $\textit{{v}}_{1}$ and $\textit{{v}}_{2}$ are two solutions of (5). Then, $\hat{\textit{{v}}}_{1}=\textit{{v}}_{1}-\Psi$ and $\hat{\textit{{v}}}_{2}=\textit{{v}}_{2}-\Psi$ satisfy (6) and so for all $\textit{{w}}\in\mathbf{W}$
@@
and since $\frac{\mathcal{C}_{NL}\|\hat{\mathbf{f}}\|_{H^{-1}(\Omega)}}{(\nu-\varepsilon)^{2}}<1$ we conclude that $\hat{\textit{{v}}}_{1}-\hat{\textit{{v}}}_{2}=0.$
@@ 2.1 Existence of very weak solution
For linear problems, like the Poisson problem [apel2016] or the Stokes problem [hamouda2017, duran2020] with Dirichlet condition in $L^{2}(\Gamma)$, the corresponding very weak solution u belongs to the space $\mathbf{L}^{2}(\Omega)$. However, in our problem if $\textit{{u}}\in\mathbf{L}^{2}(\Omega)$ to ensure that
@@
We note that, if we consider the solution of the variational problem with a regular Dirichlet data (for example, in $\mathbf{H}^{1/2}(\Gamma)$), in which case $(\textit{{u}},p)\in\mathbf{H}^{1}(\Omega)\times L^{2}_{0}(\Omega)$, then u is also a very weak solution. Indeed, from 1.4.4.5 in [Grisvard1985], we have that $W^{2,4/3}(\Omega)\subset W^{1,4}(\Omega)$, and since $L^{4}(\Omega)\subset L^{2}(\Omega)$, we can conclude that if $\phi\in\mathbf{X}$ then $\phi\in\mathbf{H}^{1}(\Omega)$. On the other hand, since $W^{1,4/3}(\Omega)\subset L^{4}(\Omega)$ (see 1.4.4.5 in [Grisvard1985] again), we can also deduce that if $q\in M$, then $q\in L^{2}_{0}(\Omega)$.
For our subsequent analysis, it is important to note that $\mathbf{H}^{1}(\Omega)\hookrightarrow\mathbf{L}^{4}(\Omega)$ [Adams1975, Grisvard1985]. Furthermore, given $\textit{{v}}\in\mathbf{H}^{1}(\Omega)$ there exists $C_{I}>0$ such that
@@
Within our analytical framework, certain regularity assumptions on the solution of the classical Stokes problem are required.
@@
$\left\{\begin{array}[]{rcc}-\nu\Delta\textit{{w}}+\nabla\pi&=\mathbf{F}&\textup{in }\,\Omega\\ \textup{div}\ \textit{{w}}&=0&\textup{in }\,\Omega\\ \textit{{w}}&=0&\textup{on }\,\Gamma\end{array}\right$.
has a unique solution $(\textit{{w}},\pi)\in\mathbf{W}^{2,4/3}(\Omega)\times W^{1,4/3}(\Omega)\setminus\mathbb{R}$ and there exists a constant $C_{S}$ such that:
$\|\textit{{w}}\|_{W^{2,4/3}(\Omega)}+\|\pi\|_{W^{1,4/3}(\Omega)}\leq C_{S}\|F\|_{L^{4/3}(\Omega)}$
@@
As stated in [marusic2000], if $\Omega$ is a bounded $C^{1,1}$ domain, the Assumption 1 is satisfied according to Theorem 3 of [amrouche-girault]. The assumption also holds if $\Omega$ is a convex polygon (see [girault1986, Theorem 5.4, Reamark 5.6] and [grisvard1979]).
From now on, we assume that $\Omega$ satisfies Assumption 1.
The first important result to consider from Marušić-Paloka work [marusic2000] is the existence and uniqueness of the very weak solution u for the case $\mathbf{f}=0$.
@@
Let $\mathbf{g}\in\mathbf{L}^{2}(\Gamma)$ satisfying (1), $\mathbf{f}=0$ and $\mathbf{L}_{\Gamma}=\{\textit{{w}}\in\mathbf{L}^{4}(\Omega),\quad\textup{div}\ \textit{{w}}=0,\quad\textit{{w}}\cdot\mathbf{n}\in L^{2}(\Gamma)\}$. If $\|\mathbf{g}\|_{L^{2}(\Gamma)}$ is small enough, then there exists a unique very weak solution $\textit{{u}}\in\mathbf{L}_{\Gamma}$ of (2). Furthermore, there exists a constant $C_{1}=C_{1}(\Omega)>0$ such that
@@
The result follows from [marusic2000, Theorem 4]. Indeed, if the Assumption 1 is satisfied, the estimate provided in [marusic2000, Lemma 2] can be obtained. Consequently, as the author note in their Remark 3, all the results from that work are applicable. ∎